1.5. Using the Morse Potential equation in the book, the C=O bond can be described with the following three parameters: D = 724 kJ/mol, a = 1.81 A., and ro-1.229 А. For this problem, you will come up with an estimate for a spring constant that can be used to model the strength of this bonding interaction. You will need to first download the Excel file PS3-Bond Energy.xlsx from Canvas and use it to do some calculations. This Excel file already has a data column (column A) for the bond length and is set up so that you can enter some necessary parameters in the table at the top and then use these parameters in equations to complete two new columns of data, one for the Morse energy and one for the Hookes Law spring energy. The two plots on the right in the Excel file will then automatically fill in, and they will update to reflect changes in your parameters in the top table. You should be familiar with using $signs in Excel formulas to lock a variable in an equation to a specific cell (or row or column) so that it does not change when you fil an equation down a column by dragging the corner of a given cell DO NOT TURN IN PRINTOUTS OF THE ENTIRE SPREADHEET OR THE PLOTS-YOU JUST NEED THE NUMERICAL ANSWERS a) In the Excel file, program in the equation for the Morse Potential (into column B) and verify that the plotted curve has the correct shape and appears to look OK. What is the energy of the bond at a bond distance of 0.5 A? b) Now program the appropriate equation for modeling the bond energy using Hookes Law into the next spreadsheet column (column C), being sure to enter the relevant parameters in the table at the top and using them in your equation. You can then manually change the parameter for the spring constant to easily change the curve for the modeled bond energy. What value of the spring constant gives a good fit to the Morse curve? (You can do this by simple eyeballing of the plot to get a reasonable answer) What is the probability of observing the bond being stretched to a length of 1.350 A? (You can use a value of 2.5 kJ/mol for kbT when doing this calculation.) c) Probability using the Morse Potential: Probability using your Hookes Law Model:If you could explain how to do it that'd be great or if you know the answers that would also be really great.
Answer #1

Hello. I can't download Excel file. I can simply calculate the value for.

1.5 The parameters a is related with spring constant

  • a={\sqrt {k_{e}/2D_{e}}},

where ke is the force constant at the minimum of the well.

a = 1.81\AA-1 and De = 724 kj/mol

Ke = a^2× 2De = 4.743 × 10^23 N/m

A) Energy at 0.5\AA

Energy V (r) = De(1-e-a(r-ro) )2

De = 724 kJ/mol, a = 1.81 A-1, r = 0.5A, ro = 1.229A

V(r) = 724 × 1000 × ( 1- exp(1.81 (0.5-1.229)))^2

V(r) = 5441.50 kJ/mol

I can't not calculate further